The perimeter of an isosceles triangle is 28cm. What is the maximum area possible? 


1
Expert's answer
2021-10-05T15:09:39-0400

Let b be the length of the base, and a the length of each of the two equal legs.

Triangle area:


"S=\frac{b\sqrt{a^2-b^2/4}}{2}"


"b=28-2a"


"S=(14-a)\sqrt{28a-196}"


"S'(a)=-\sqrt{28a-196}+\frac{14(14-a)}{\sqrt{28a-196}}=0"


"-28a+196+196-14a=0"


"a=28\/3" cm

"b=28-56\/3=28\/3"


"S_{max}=\frac{28\sqrt{784/9-784/36}}{6}=\frac{196\sqrt{3}}{9}"